OpenStudy (anonymous):

can anyone help me find 't' when v=50 in the following equation please? (including steps) v=100sin(200*pi*t+(pi/4)) Thanks in advance!

Vocaloid (vocaloid):

v=100sin(200*pi*t+(pi/4)) substitute 50 for v, giving us 50=100sin(200*pi*t+(pi/4)) divide both sides by 100, giving us (1/2) = sin(200*pi*t+(pi/4)) with me so far?

Vocaloid (vocaloid):

@MathBiff

Vocaloid (vocaloid):

next step is to take the arcsin of both sides, which will "undo" the sin on the right side arcsin(1/2) = 200*pi*t+(pi/4) re-arrange to solve for t

OpenStudy (anonymous):

cheers @Vocaloid. That's more or less where i got up to, the next step would be to subtract (pi/4) from both sides the divide both sides by 200*pi right? (arcsin(1/2)-(pi/4))/200*pi=t (30-(pi/4))/628.3=t The other examples which I found on this question had 30+(pi/4). Just trying to find out which is right.... Here's the other example: When V=50V then 50V= 100Sin(200*pi*t+0.785) 50/100=Sin(200*pi*t+0.785) (200*pi*t+0.785)=arcsin 50/100 =30degrees or 0.524rad 200*pi*t= 0.524+0.785 = 1.309 So when V=50v Time t = 1.309/200*pi = 2.08ms

Vocaloid (vocaloid):

yeah, your answer looks right to me, I'm not sure how they got 30+(pi/4)

OpenStudy (anonymous):

Ah cool. Thanks for that. I'm not exactly a math wizard so wasn't sure if I'd missed a trick somewhere. Have a good day!